OpenStudy (marigirl):

Shane gets $1.50 more per hour than Ben. Shane works for 8 hours and Ben for 9 hours and together they make a total of $182. what is the hourly rate for each?

OpenStudy (anonymous):

8.6 and 10.1?

OpenStudy (anonymous):

Let x = the amount of money Ben makes per hour Let y = the amount of money Shane makes per hour So we know that together they make $182.00. That would look like this: x+y = 182 We also know that Shane makes $1.50 more than Ben, each hour, therefore: y=x+1.50 Now that you've got 2 equations, can you take it from there?

OpenStudy (anonymous):

you forgot to add the amount of hours they work?

OpenStudy (anonymous):

include*

OpenStudy (anonymous):

Oh! You're right! Look at that! :) My bad!

OpenStudy (anonymous):

So the first equation I wrote : x+y=182 It should instead be: 9x+8y=182 I'm sorry about that.

OpenStudy (marigirl):

so that would be 8x+9y=182 and x+1.50=y

OpenStudy (marigirl):

awesome! thanks you soo much

OpenStudy (anonymous):

i feel so dumb, i don't know how to do these problems anymore -_-

OpenStudy (anonymous):

That's right! :) Good! You're Welcome!!

OpenStudy (marigirl):

your not dumb, you are just simply expanding your knowledge. like me.

OpenStudy (anonymous):

Yeah, don't feel bad. I do these all day and even I messed that one up. :)

OpenStudy (anonymous):

wait so what's the ansewr? for my peace of mind

OpenStudy (anonymous):

What did you get Marigirl?

OpenStudy (marigirl):

Shane gets $11.50 and Ben $10. Thanks so much guys.!

OpenStudy (anonymous):

Yay! That's what I got too! :) You're quite welcome!